OpenStudy (anonymous):

can someone please explain to me how to use factoring in solving quadratic equations and possibly provide an example so I can see the steps?

OpenStudy (saifoo.khan):

Heya!

OpenStudy (anonymous):

you should see in each equation what is comman and get that as a comman element (factor) for example 3x+6=0 for factoring this you may see what is comman , we can see that 3 is comman and it will be the factoring: 3(x+2)

OpenStudy (anonymous):

HI Saif!!!!!!

OpenStudy (anonymous):

let us take the example 3x² - 5x -8 = 0 let us write the general form ax² + bx + c = 0 so we write given equation as 3x² + (-5)x + (-8) = 0 first we shud find product of a and c, here a=3 and c = -8 so a*c = 3 * (-8) = -24

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

can you explain it in words as well Harkiat?

OpenStudy (anonymous):

pls tell me which part is unclear??

OpenStudy (anonymous):

from the given quadratic equation, we have to first locate values of a and c and find their product...

OpenStudy (anonymous):

so using factoring is finding a common number?

OpenStudy (anonymous):

no, it is finding factors of the product ac so that their sum is equal to b, where a is coefficient of x² b is coefficient of x and c is the constant

OpenStudy (anonymous):

do u hv a SKYPE or YAHOO MESSANGER id where we can have voice chat it will be easier and quicker that way....

OpenStudy (anonymous):

i do have yahoo messenger, but i don't know if voice will work on this computer but we can try

OpenStudy (anonymous):

i have both on my home computer

OpenStudy (anonymous):

downloading SKYPE will hardly take 5 minutes...

OpenStudy (anonymous):

or you can go into your yahoo email acoount and yahoo msngr is availbl there too

OpenStudy (anonymous):

my yahoo msngr id is hakki.mesho

OpenStudy (anonymous):

i am on a work computer I can't do that. but I can use yahoo messenger

OpenStudy (anonymous):

I am trying to figure out how to add you on this new yahoo mail thing.

OpenStudy (anonymous):

can u use voice chat in yahoo msnger???

OpenStudy (anonymous):

i think so.

OpenStudy (anonymous):

i don't know i can't fgigure the one out from my email. let me go to my computer one it does more.

OpenStudy (anonymous):

yea but it doesn't give me an option to add anymore try to add me rachel_manring

OpenStudy (anonymous):

i figured it out but it won't let me add you

OpenStudy (anonymous):

rachel, still there????